OpenStudy (anonymous):

For the simple harmonic motion equation D=4 sin(8pit) , what is the period? If necessary, use the slash (/) to represent a fraction.

OpenStudy (anonymous):

What have you tried so far?

OpenStudy (anonymous):

i think is 1/2 but i'm not sure

OpenStudy (anonymous):

i mean 1/4

OpenStudy (anonymous):

and you are right

OpenStudy (anonymous):

so is 1/4?

OpenStudy (anonymous):

y

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks ;)

OpenStudy (anonymous):

Yup. Look at what is in parenthesis for sin 8pi*t set equal to 2pi and solve for t 8pi*t = 2pi t = (2pi)/(8pi) = (1/4)

OpenStudy (anonymous):

More generally if period of \( f(x) \) be \(T\) then \( kf(ax+b)\) will have a period of \(\huge \frac T {|a|} \)